8 g NaOH H2SO4 are present in 1 litre of the solution.What is its PH

Grade:

1 Answers

Himanshu Rathour
5 years ago
Moles of NaOH=8/40
=0.2moles
Moles of H2SO4=8/98
≈0.08moles
Now NaOH and H2SO4 are both strong base and strong acids thus they are completely ionised
Thus
Moles of OH¯=0.2
Moles of H+=0.08
Hence net solution will be basic
Moles of OH¯=0.12
pOH=-log(0.12)
=0.92
pH=14-0.92
pH=13.08